OpenStudy (anonymous):

another one 22-2x+3=38x+10-3x

OpenStudy (anonymous):

x = 15/37

OpenStudy (anonymous):

thanks, I came up with 18

OpenStudy (anonymous):

22 + 3 - 10 = 38x - 3x +2x 15 = 37 x x = 15/37

OpenStudy (anonymous):

thanks